OpenStudy (anonymous):

16) How many photons are contained in a burst of yellow light (589 nm) from a sodium lamp that contains 609 kJ of energy? A) 3.37 × 1019 photons B) 3.06 × 1030 photons C) 1.81 × 1024 photons D) 4.03 × 1028 photons E) 2.48 × 1025 photons

OpenStudy (anonymous):

Okay, I am having trouble getting the right answer. I am getting 1.81 x 10 ^ 16

OpenStudy (aaronq):

find the energy of one photon: \[E=h \frac{ c }{\lambda }\] divide the energy given by the energy of one photon of that wavelength

OpenStudy (anonymous):

What I've done so far is convert wave length to m and energy to j. E photon = h * x / wave length E = (6.626 x 10^-43)(3.00 x 10^8) / 587 ^ -9 = 3.38 x 10 ^18 J 3.38 x 10 ^18 J x 1000 kj / 1 j = 3.37 x 10 ^ 16 Kj 609 kJ/ 3.37 x 10 ^ 16 Kj = 1.81 x 10 ^ 16

OpenStudy (anonymous):

E = (6.626 x 10^-34)(3.00 x 10^8) / 587 ^ -9 = 3.38 x 10 ^19 J 3.38 x 10 ^19 J x 1000 kj / 1 j = 3.37 x 10 ^ -16 Kj 609 kJ/ 3.37 x 10 ^ 16 Kj = 1.81 x 10 ^ 18 but the answer is 1.81 × 10^24 photons

OpenStudy (anonymous):

@Preetha Hello, can you help me figure out what I'm dong wrong here?

OpenStudy (aaronq):

3.38 x 10 ^-19 J should be negative

OpenStudy (aaronq):

then 3.38 x 10 ^18 J x 1kJ/1000 J

OpenStudy (aaronq):

you're converting from J to kJ.. just like meters to kilometres, you wouldn't multiply you would divide

OpenStudy (anonymous):

oh i see. got it!

OpenStudy (aaronq):

good stuff

OpenStudy (anonymous):

I got j and kJ mixed. Usually killo means 1000 o_O

OpenStudy (aaronq):

kilo is 1000.. there are 1000J in 1kJ

OpenStudy (anonymous):

Makes sense. Thanks! This cleared my problem.

OpenStudy (aaronq):

awesome, no prob dude !